OpenStudy (anonymous):

Find the discrinmant of 3x^2 - 7x + 2 = 0 and what does discrimnant mean

OpenStudy (anonymous):

Please help

Parth (parthkohli):

Discriminant in \(ax^2 + bx + c = 0\) is \(b^2 - 4ac\)

Parth (parthkohli):

Here the discriminant would be \((-7)^2 - 4(3)(2)\)

OpenStudy (anonymous):

25?

OpenStudy (anonymous):

what does this word mean?

Parth (parthkohli):

Yeah. Discriminant determines the number of solutions in a quadratic equation :)

Parth (parthkohli):

If the discriminant is a perfect square, then we have two integer solutions.

OpenStudy (anonymous):

what does it mean when its 25

OpenStudy (unklerhaukus):

For a quadratic equation of the form \[ax^2+bx+c=0\] The discriminant is some times written like this\[\Delta=b^2-4ac\]

OpenStudy (aravindg):

well the discriminant is a no that gives us the idea how the roots of the quadratic equation will be

OpenStudy (unklerhaukus):

the quadratic formula used to finds solutions of x in a quadratic equation \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} =\frac{-b\pm\sqrt\Delta}{2a}\]

OpenStudy (aravindg):

if the discriminant is irrational the roots are irrational conjugates , if the discriminant is 0 there is only 1 root , if discriminant >0 there are 2 real and distinct roots , if discriminant <0 roots are imaginary hope that clears ur doubt

OpenStudy (unklerhaukus):

the discriminant is used to determine the number of solutions for x ie the number of times the quadratic curve cuts the x-axis \(\Delta >0 \) if the discriminant is a positive number there are 2 solutions, ie the parabola cuts the x-axis twice \(\Delta =0 \) if the discriminant is zero there is one solution ie the parabola touches the x-axis at one point \(\Delta <0 \) if the discriminant is a negative number there are no real solutions ie the parabola does not cut the x-axis

OpenStudy (unklerhaukus):

if the discriminant is a positive number that is also a square number the solutions/roots of the equation are rational numbers the points where the parabola cuts the x-axis are rational point on the x-axis

OpenStudy (unklerhaukus):

for example \[4x^2 - 3x -1 = 0\] \[\Delta=b^2-4ac=(-3)^2-4\times(4)\times(-1)\]\[\qquad\qquad\qquad=9+16\]\[\qquad\qquad\qquad=25\] \(\Delta>0\) so there are two solutions \(25\) is also a square number so the roots are rational the quadratic formula \[x=\frac{-(-3)\pm\sqrt {25}}{2\times-3}=\frac{3\pm5}{-6}\] \[x_{1,2}=-\frac {4}{3},\frac 13\] which are two rational roots